OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 10, x = 3. Find y when x = 5. help me please

OpenStudy (anonymous):

The equation is y = k(2x) If y = 10, x = 3, then plug these 2 numbers in to the equation to find k. 10 = k(6) k = 10/6 Now we know k = 10/6 or 5/3, so our equation is y = (5/3)(2x) Lets now answer the question...find y, when x = 5, so y = (5/3) (10) y = 50/3.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

welcome.